OpenStudy (anonymous):

x-2 OVER x-1= x+2 OVER x+4

Directrix (directrix):

(x - 2)/(x - 1) = (x+2)/(x+4) x^2 + 2x - 8 = x^2 + x -2 2x -8 = x - 2 x = 6 ----.> Test x =6 in each denominator of the original equation to ensure that 6 does not create a zero and cause the expression to become undefined. Will x = 6 cause (x-1) = 0? No Will x = 6 cause (x+4) = 0? No. Therefore, the answer is x = 6.